# A comprehensive evaluation of spatial co-execution on GPUs using MPS and MIG technologies

### 24 Apr 2026 v1

@@ Abstract
To mitigate the increasingly common underutilization of computational resources in modern GPUs, spatial sharing methods enable multiple applications to use them simultaneously. This work presents a comprehensive evaluation of NVIDIA’s primary technologies to achieve that goal: Multi-Process Service (MPS) and Multi-Instance GPU (MIG). Our findings reveal a crucial trade-off between MPS’s flexibility and MIG’s isolation, and provide many key insights for improving the co-execution strategy according to job profiles. In the most favorable scenarios, MPS improves performance by up to 30% and reduces energy by about 20%, using its provisioning option to avoid resource monopolization. However, under memory contention, it suffers severe degradation, worsening performance by around 30%. Conversely, MIG’s full hardware isolation resolves memory contention, leading to more consistent improvements, but these gains are tempered by higher overhead, and its rigid scheme can degrade performance in certain cases.
@@ 1 Introduction
For spatial resource sharing, NVIDIA, the prominent leader in the GPU market, offers the Multi-Process Service (MPS) [mps] and Multi-Instance GPU (MIG) [mig] technologies, whose main difference lies in their approach towards isolation. Whereas MPS shares resources among applications—with an option to limit the maximum percentage of SMs, but not to restrict memory system usage—MIG is based on the physical partitioning of the device into smaller, independent instances, each with its own fully isolated compute and memory resources. Consequently, MPS provides flexible sharing but may suffer from memory-system contention that can violate QoS (Quality of Service) or SLA (Service Level Agreement), while MIG provides a stronger isolation at the cost of coarser execution patterns and more overhead. However, these trade-offs have been insufficiently evaluated in the prior literature [Li2022_MISO, Zhang2024, Zhao2022], which lacks comprehensive comparisons between both technologies, and especially, in-depth analyses and characterization with diverse workload combinations.
This work contributes to this area by performing a multi-factorial evaluation of different co-execution schemes on NVIDIA GPUs. Our main contributions are as follows:
@@ 2.2 Time-slicing
The primary benefit of this time-slicing is fairness among processes, which can particularly reduce the average completion time in multi-application environments. However, it can also yield utilization benefits by eliminating idle periods during operations such as host-to-device (or vice versa) transfers, CPU computations between kernels, or I/O waits. As we will show, our experiments indicate that the aggregate performance of the applications improves slightly with this technique, suggesting that its benefits marginally outweigh the overhead from context switches. While it is possible to modify the scheduling policy and the duration of time slices using tools like Kubernetes with the NVIDIA GPU Operator (e.g., equal-share, fixed-share, long/medium/short slices), we will use the default policy simply as a baseline against which to compare the more advanced methods described below. These methods will be more ambitious than time-slicing, achieving superior improvements in performance and energy consumption, but they require explicit configuration actions that can sometimes be complex (involving scheduling decisions) and incur higher overhead.
@@ 3.2 Co-execution schemes
  • MPS-best-prov: Spatial sharing with MPS, but provisioning the maximum percentage of resources that each application can use. This provides a degree of isolation when the total allocation is under 100% (see Section 2.3). We select the optimal provisioning (-best-prov) by sweeping through resource distributions in 10% increments, choosing the split that minimizes the makespan (i.e., the completion time of the last application to finish).
@@ 3.3 Evaluation metrics
This work will analyze energy consumption by examining the fundamental trade-off between performance and power draw, dissecting the independent contribution of each factor. While higher performance reduces execution time, it often requires a power increase. In this context, we measured the total Joules consumed during the execution with the nvmlDeviceGetTotalEnergyConsumption function from the NVML library [nvidia_nvml]. This value is divided by the total execution time to calculate the power (in watts), which will be important in the analysis, and is also presented in relation to the values measured for the applications running alone.
To deepen our analysis, we complement these primary metrics with resource utilization data. We employ Nsight Compute (NCU) to obtain fine-grained throughput metrics—specifically, peak bandwidth utilization for memory and peak SM performance per kernel—in select scenarios (Footnote 3 and Table 4). However, NCU is generally incompatible with MPS due to the unified context, which prevents the accurate attribution of hardware counters to individual processes. Therefore, following prior studies on MPS [Elvinger2025, Robroek2024, Weaver2024, Zhao2023MuxFlow], we primarily rely on the DCGM tool [NVIDIA_DCGM] to track memory and Compute (SM) utilization. DCGM performs periodic sampling (every 200 ms in our case), dividing each period into 100 equal-time slots for which it checks resource activity to obtain a utilization percentage. Finally, these percentages are averaged over time. In addition, to quantify fairness by assessing the slowdown disparity between co-executed applications (Section 4.2), we employ the Gini coefficient as a statistical measure of inequality [giudici2025RGA, giudici2025measuring].
@@ 3.4 Workloads
KMeans Test 11 different GPU implementations of the popular clustering algorithm. $131072$ points $32$ dimensions $32$ centroids 2
@@
Enformer [Avsec2021Enformer] Hybrid CNN-transformer for predicting gene expression from large DNA sequences. It has $\sim 230$M parameters. DNA seq. size 196196k Batch size 1 50
PosEGNN [ibm_pos-egnn_2023] Graph Neural Network of IBM for predicting molecular potential energy and inter-atomic forces. $\sim 6$M parameters. Dynamic simulation for $80$ atoms 150
Regarding the workloads, we selected various tests from the Altis GPU benchmark suite [Altis], which has been widely used in other evaluations of the literature [DeLaCalle2025, Sainz2022, VillarrubiaDRL2026], as well as four popular deep learning models used for inference (particularly one applied to genomics and another to materials analysis). Most of the benchmarks are configured by a level parameter from 1 to 5, that adjusts the input configuration to vary the computational demand (level=1level\!=\!1 uses the least demanding configuration and level=5level\!=\!5 the most). In other cases, that configuration can be determined manually. Table 1 lists the chosen applications (with a brief description of each), the level or problem size used, and the number of passes set for each test. Correspondingly, Table 2 shows their standalone execution times in an NVIDIA A30, along with SM and memory utilization metrics (average and maximum), and their energy and power consumption with full GPU access. It should be noted that for the benchmarks and workloads used in this study, the program that launches the test automatically selects the thread configuration (number of blocks and threads per block for deployment on the GPU as explained in Section 2.1) based on the problem size and the GPU’s characteristics.
@@
RDL Enformer 5.98 81.23 100 44.70 67 599.21 100.20
RDL PosEGNN 5.11 34.98 60 12.47 18 333.42 65.25
As can be observed, the average SM utilization is quite low (below 40%) in most cases, which indicates sparse resource usage, making them good candidates for co-execution. The exception among the Altis applications is GeMM, with 78% average SM utilization. While this might suggest it is not a good candidate for co-execution, its memory utilization is very low (7.87% on average), leaving memory resources free for other applications. Furthermore, the Enformer model demonstrates high average SM utilization (81.23%), combined with moderate memory utilization (44.70%). This is a common workload with a substantially different profile from the others, which will provide revealing conclusions.
To ensure the results were illustrative, we sought to balance the application runtimes without altering their inherent nature. This is because if one application finishes much later than another, the latter spends a significant amount of time running alone, which dilutes the co-execution results. This effect is particularly pronounced with approaches like MPS, where the remaining application can expand to use all GPU resources once the other has completed. While one can imagine applications in a continuous co-execution stream, for practical purposes, we must evaluate them over a finite horizon. Thus, the number of execution passes was adjusted so that the test durations were similar; Table 1 specifies the passes for each, while Table 2 shows that the execution times are all within a narrow range of 5 to 6 seconds.
@@
Although the Altis tests correspond to real applications, they are designed to test the GPU in a particular way, and limiting our study to them could restrict its validity. To broaden the workload spectrum, we incorporated four prominent deep learning applications, all executed within the PyTorch framework. For standard inference tasks, we utilized ResNet50 and BERT models via the TorchVision library. Additionally, to represent scientific domain-specific workloads, we included the Enformer and Pos-EGNN neural networks—targeting genomic prediction and materials analysis, respectively—sourced from the Hugging Face repository. Their evaluation datasets and batch sizes are specified in Table 1. We classify these models into a new category, Type RDL (Real Deep Learning), which we will be used for additional validation. As shown in Table 2, the selected models exhibit diverse resource usage profiles: BERT demonstrates moderate SM utilization with limited memory pressure; Enformer is resource-intensive in both dimensions; and ResNet50 or Pos-EGNN impose significantly lighter loads. These varied profiles help corroborate the findings from the Altis benchmark tests.
@@ 4.1 Impact of memory resource contention
Figure 6 compares the individual slowdown of the co-execution schemes for each pair of M-type applications. For reference, the black horizontal line marks a slowdown of $2\times$, which can be considered a satisfactory reference, since it means that the application runs concurrently with another without either exceeding twice its solo execution time. It is particularly striking that MPS-based schemes perform so poorly for these combinations: they systematically degrade both applications’ performance, reaching slowdowns of up to 2.5 (especially for pairs including Sort). The underlying cause is revealed in Figure 7, which illustrates a considerable surge in memory utilization with MPS compared to time-slicing or MIG, evident in both average and peak values. In cases of greater slowdown (Sort test), it can be observed that MPS exhibits an average memory utilization of roughly 60%, rising to nearly 80% for the SortSort pair (the highest slowdown), with peaks of 100% for applications such as GUPS. By contrast, time-slicing barely exceeds 20%—values that closely match the solo-execution metrics in Table 2—while executions under MIG show only a slight increase over those solo values. This reflects prolonged DRAM engagement, which points to contention in the GPU memory subsystem shared under MPS. The likely causes are an increased rate of misses in lower-level caches and simultaneous access attempts saturating the L2 cache and DRAM bandwidth [^1] As previously mentioned, finer-grained profiling, such as that offered by NCU tool, is unfortunately incompatible with MPS, preventing analysis of lower-level metrics like cache misses.. In the rest of the executions, the differences in memory usage are not as exaggerated, but still very noticeable, which justifies the performance degradation observed in the slowdown.
@@
Although the co-execution performance under MPS-best-prov in Figure 6 remains poor, it represents a slight improvement over basic MPS. Recall that if the sum of the SM percentages does not exceed 100% (as in this case in MPS-best-prov as explained in Section 3.2), isolation is achieved at the SM level. This means no single SM hosts threads from both applications, which prevents contention from sharing resources like the L1 cache. While this could potentially be counterproductive for warp interleaving, the thread and block configurations for each application are generally sufficient to achieve 100% theoretical occupancy (as verified with NCU tool for numerous kernels). Figure 7 shows a slight decrease in memory utilization that hints at this improvement, yet this metric is still much higher than that of time-slicing or MIG, since SM provisioning does not isolate the rest of the memory subsystem as these workloads demand.
@@
Table 3: Duration (in seconds) and scaling efficiency according to the number of slices $s$ of the MIG instance ($E=t(1)/(t(s)\cdot s)$).
@@ 4.2 Resource monopolization and fair sharing
A striking observation from Figure 10 is the asymmetric performance degradation on pairs involving the GeMM benchmark. With MPS, the application co-executing with GeMM suffers significant slowdown, often exceeding $2\times$ (see GeMM paired with GUPS or SRAD). Conversely, with MIG, it is GeMM itself that is penalized, while the co-running application remains largely unaffected. Quantifying this inequality, the average Gini coefficient for pairs involving GeMM reaches 0.35 for MPS and 0.12 for MIG, compared to just 0.03 for non-GeMM pairs, and dropping to 0.01 for these same GeMM pairs under Time-slicing and MPS-prov-best. We attribute this behavior to the distinct resource requirements of GeMM, which, as detailed in Section 3.4, is significantly more compute-intensive than the other workloads. Consequently, under MPS, GeMM monopolizes the SMs, effectively starving the co-running application even if the memory system remains underutilized. In contrast, under MIG, GeMM is strictly confined to a single instance (half the GPU resources in a 2–2 partition). This isolation prevents it from accessing the necessary compute resources, causing its performance to degrade by more than $2\times$, whereas the less demanding co-running application operates efficiently within its isolated slice.
@@
Figure 11: Co-execution slowdowns for pairs of RDL and M-type workloads (the best MPS provisioning is indicated).
@@ 4.3 Impact of partitioning granularity and flexibility
As observed in Figures 10 and 11, GeMM, ResNet50 and Enformer suffer from significant performance degradation when co-executing on 2-slice instances (much more than double the solo runtime), that presumably would not occur with 3 slices (if such an instance were possible). In these cases, it could be argued that co-locating applications is not as critical, as the problem arises from their already high resource utilization (see Table 2). This means MIG partitioning lacks the granularity and flexibility to accommodate such fine-grained adjustments.
@@
It is worth noting several nuances regarding this issue. First, we have illustrated this on an A30 GPU because its partitioning space is more limited: it only has 4 slices, 3 instance sizes (1, 2, and 4), and 5 possible partitions, making it unfeasible, for example, to assign 3 or 4 slices to the application with higher memory demand and 2 to the other. On bigger GPUs such as the A100, H100, or B200, which feature higher throughput and greater MIG granularity (7 slices, 5 possible instance sizes, and up to 19 partitions), this issue becomes less relevant because it is easier for applications to adapt well to one of the available device partition proportions. Second, MIG includes Compute Instances (see Section 2.4), which allow memory to be shared among concurrent applications within the same instance without sharing the SMs. With these instances, it is possible to obtain results similar to those observed with MPS, although in general cases, this implies predicting and controlling interference to avoid contention problems, which is not always easy, as some scheduling proposals have noted [Xiao2018, Zhang2024]. Finally, the ideal solution would be a MIG-like, but asymmetric, isolation: for example, being able to form an instance with 2 compute slices and 4 memory slices, all of them fully isolated. However, that seems infeasible or technically very difficult, at least in the short term; currently, MIG uses isolation at the GPC level, but it would likely require the ability to isolate resources at a lower, decoupled level.

### 29 Apr 2026 v2

@@ Abstract
To mitigate the increasingly common underutilization of computational resources in modern GPUs, spatial sharing methods enable multiple applications to use them simultaneously. This work presents a comprehensive evaluation of NVIDIA’s primary technologies to achieve that goal: Multi-Process Service (MPS) and Multi-Instance GPU (MIG). Our findings reveal a crucial trade-off between MPS’s flexibility and MIG’s isolation, and reveals many key insights for improving the co-execution strategy according to jobs profiles. In the most favorable scenarios, MPS improves performance by up to 30% and reduces energy by about 20%, using its provisioning option to avoid resource monopolization. However, under memory contention, it suffers severe degradation, worsening performance by around 30%. Conversely, MIG’s full hardware isolation resolves memory contention, leading to more consistent improvements, but these gains are tempered by higher overhead, and its rigid scheme can degrade performance in certain cases.
@@
This is a preprint of an article published in The Journal of Supercomputing.
The final version is available at https://doi.org/10.1007/s11227-026-08312-z
@@ 1 Introduction
For spatial resource sharing, NVIDIA offers the Multi-Process Service (MPS) [mps] and Multi-Instance GPU (MIG) [mig] technologies, whose main difference lies in their approach towards isolation. Whereas MPS shares resources among applications —with an option to limit the maximum percentage of SMs, but not to restrict memory system usage— MIG is based on the physical partitioning of the device into smaller, independent instances, each with its own fully isolated compute and memory resources. Consequently, MPS provides flexible sharing but may suffer from memory-system contention that can violate QoS (Quality of Service) or SLA (Service Level Agreement), while MIG provides a stronger isolation at the cost of coarser execution patterns and more overhead. However, these trade-offs have been insufficiently evaluated in the prior literature [Li2022_MISO, Zhang2024, Zhao2022], which lacks comprehensive comparisons between both technologies, and especially, in-depth analyses and characterization with diverse workload combinations.
This work contributes to this area by performing a multi-factorial evaluation of different co-execution schemes on modern GPUs. Our main contributions are as follows:
@@ 2.2 Time-slicing
The primary benefit of this time-slicing is fairness among processes, which can particularly reduce the average completion time in multi-application environments. However, it can also yield utilization benefits by eliminating idle periods during operations such as host-to-device (or vice versa) transfers, CPU computations between kernels, or I/O waits. As we will show, our experiments indicate that the aggregate performance of the applications improves slightly with this technique, suggesting that its benefits marginally outweigh the overhead from context switches. While it is possible to modify the scheduling policy and the duration of time slices using tools like Kubernetes with the NVIDIA GPU Operator (e.g., equal-share, fixed-share, long/medium/short slices), we will use the default policy simply as a baseline against which to compare the more advanced methods described below. These methods will be more ambitious than time-slicing, achieving superior improvements in performance and energy consumption, but they require explicit configuration actions that can sometimes be complex (involving scheduling decisions) and incur in a higher overhead.
@@ 3.2 Co-execution schemes
  • MPS-best-aprov: Spatial sharing with MPS, but provisioning the maximum percentage of resources that each application can use. This provides a degree of isolation when the total allocation is under 100% (see Section 2.3). We select the optimal provisioning (-best-aprov) by sweeping through resource distributions in 10% increments, choosing the split that minimizes the makespan (i.e., the completion time of the last application to finish).
@@ 3.3 Evaluation metrics
This work will analyze energy consumption by examining the fundamental trade-off between performance and power draw, dissecting the independent contribution of each factor. While higher performance reduces execution time, it often requires a power increase. In this context, we measured the total Joules consumed during the execution with the nvmlDeviceGetTotalEnergyConsumption function from the NVML library [nvidia_nvml]. This value is also divided by the total execution time to calculate the power (in watts), which will be important in the analysis, and is also presented in relation to the values measured during the standalone execution of the applications.
Finally, to understand and analyze the results, these primary metrics are supplemented by auxiliary metrics on resource utilization. One option is the throughput metrics offered by Nsight Compute (NCU), which include peak bandwidth utilization percentages for different memory levels and the achieved percentage of peak SM performance for each kernel (it will be used in Footnote 3 and Table 4). However, NCU is not compatible with the MPS service because the unified MPS context creates ambiguity in how the tool has to take and organize hardware counters, so we cannot use it in most of our experiments (though we will use it in a few cases where it provides particularly illustrative insights). As other studies on MPS [Elvinger2025, Robroek2024, Weaver2024, Zhao2023MuxFlow], we will instead use the memory and compute (SM) utilization percentages provided by the DCGM tool for this purpose [NVIDIA_DCGM]. This tool performs periodic measurements (every 200 ms in our case) to calculate resource utilization. Each period is divided into 100 slots, and by checking for resource activity in each slot, a utilization percentage is obtained. Finally, these percentages are averaged over time.
@@ 3.4 Workloads
KMeans Test 11 different GPU implementations of the popular clustering algorithm. 131072131072 points $32$ dimensions $32$ centroids 2
@@
Regarding the workloads, we selected various tests from the Altis GPU benchmark suite [Altis], which has been widely used in other evaluations of the literature [DeLaCalle2025, Sainz2022, VillarrubiaDRL2026]. Most of them are configured by a level parameter from 1 to 5, that adjusts the input configuration to vary the computational demand (level=1level\!=\!1 uses the least demanding configuration and level=5level\!=\!5 the most). In other cases, that configuration can be determined manually. Table 1 lists the chosen applications (with a brief description of each), the level or problem size used, and the number of passes set for each test. Correspondingly, Table 2 shows their standalone execution times in an NVIDIA A30, along with SM and memory utilization metrics (average and maximum), and their energy and power consumption when run alone with full GPU access. It should be noted that for the benchmarks and workloads used in this study, the program that launches the test automatically selects the thread configuration (number of blocks and threads per block for deployment on the GPU as explained in Section 2.1) based on the problem size and the GPU’s characteristics.
@@
As can be observed, the average SM utilization is quite low (below 40%) in most cases, which indicates sparse resource usage, making them good candidates for co-execution. The exception among the Altis applications is GeMM, with 78% average SM utilization. While this might suggest it is not a good candidate for co-execution, its memory utilization is very low (7.87% on average), leaving memory resources free for other applications. We wanted to include such an application because it represents a different profile and, as we will see, its results will be quite interesting.
To ensure the results of some experiments were illustrative, we sought to balance the application runtimes without altering their inherent nature. This is because if one application finishes much later than another, it spends a significant amount of time running alone, which dilutes the co-execution results. This effect is particularly pronounced with approaches like MPS, where the remaining application can expand to use all GPU resources once the other has completed. While one can imagine applications in a continuous co-execution stream, for practical purposes, we must evaluate them over a finite horizon. Thus, the number of execution passes was adjusted so that the test durations were similar; Table 1 specifies the passes for each, while Table 2 shows that the execution times are all within a narrow range of 5 to 6 seconds.
@@
Although the Altis tests correspond to real applications, they are designed to test the GPU in a particular way, and limiting our study to them could restrict its validity. In this context, we have also included two widely used Deep Learning applications: inference with ResNet50 and BERT models, using the TorchVision library within the PyTorch framework. Their evaluation datasets and batch sizes are specified in Table 1. We classify these models into a new category, Type RDL (Real Deep Learning), which we will use primarily for additional validation. As shown in Table 2, BERT makes moderate use of SMs and low use of memory, whereas ResNet50 is quite intensive in both. These distinct behaviors will serve to corroborate the observations obtained from the Altis benchmark tests.
@@ 4.1 Impact of memory resource contention
Figure 6 compares the individual slowdown of the co-execution schemes for each pair of M-type applications. For reference, the black horizontal line marks a slowdown of $2\times$, which can be considered a satisfactory reference, since it means that the application runs concurrently with another without either exceeding twice its solo execution time. It is particularly striking that MPS-based schemes perform so poorly for these combinations: they systematically degrade both applications’ performance, reaching slowdowns of up to 2.5 (especially for pairs including Sort). The underlying cause is revealed in Figure 7, which illustrates a considerable surge in memory utilization with MPS compared to time-slicing or MIG, evident in both average and peak values. In cases of greater slowdown (Sort test), it can be observed that MPS exhibits an average memory utilization of roughly 60%, rising to nearly 80% for the SortSort pair (the highest slowdown), with sustained 100% peaks. By contrast, time-slicing barely exceeds 20%—values that closely match the solo-execution metrics in Table 2—while executions under MIG show only a slight increase over those solo values. This reflects prolonged DRAM engagement, which points to contention in the GPU memory subsystem shared under MPS. The likely causes are an increased rate of misses in lower-level caches and simultaneous access attempts saturating the L2 cache and DRAM bandwidth [^1] As previously mentioned, finer-grained profiling, such as that offered by NCU tool, is unfortunately incompatible with MPS, preventing analysis of lower-level metrics like cache misses.. In the rest of the executions, the differences in memory usage are not as exaggerated, but still very noticeable, which justifies the performance degradation observed in the slowdown.
@@
Although the co-execution performance under MPS-best-prov in Figure 6 remains poor, it represents a slight improvement over basic MPS. Recall that if the sum of the SM percentages does not exceed 100% (as in this case in MPS-best-prov as explained in Section 3.2), isolation is achieved at the SM level. This means no single SM hosts threads from both applications, which prevents contention for shared resources like the L1 cache. While this could potentially be counterproductive for warp interleaving, the thread and block configurations for each application are generally sufficient to achieve 100% theoretical occupancy (as verified with NCU tool for numerous kernels). Figure 7 shows a slight decrease in memory utilization that hints at this improvement, yet this metric is still much higher than that of time-slicing or MIG, since SM provisioning does not isolate the rest of the memory subsystem as these workloads demand.
@@
Table 3: Duration (in seconds) and scaling efficiency according to the number of slices $s$ of the MIG instance ($E(s)=t(1)/(t(s)\cdot s)$).
@@ 4.2 Resource monopolization and fair sharing
A striking observation from Figure 10 is the poor performance of both MPS and MIG for one of the two applications in pairs involving the GeMM benchmark. With MPS, the application co-executing with GeMM is the one that experiences a significant slowdown, sometimes exceeding a factor of 2x (e.g., GeMM with GUPS or SRAD). Conversely, with MIG, it is GeMM itself that is penalized, while the co-running application is less affected. We attribute this to the fact that, as explained in Section 3.4, GeMM is a much more demanding application than the others, although mainly in terms of calculation rather than memory access. Consequently, under MPS, GeMM monopolizes the SMs, preventing the co-running application from accessing them, even if that application could better utilize the underused memory system. With MIG, GeMM is confined to an instance with half the GPU resources (one instance in a 2–2 partition), which causes performance to degrade by more than double because its demand is higher. Meanwhile, the less demanding co-running application, isolated in its own 2-slice instance, is not similarly impacted (the amount of resources is sufficient for it).
@@
Figure 11: Co-execution slowdowns for pairs of RDL and M type workloads (the best MPS provisioning is indicated).
@@ 4.3 Impact of partitioning granularity and flexibility
As observed in Figures 10 and 11, the GeMM and ResNet50 applications suffer from significant performance degradation when co-executing on 2-slice instances (more than double the runtime). This issue would presumably not occur on a 3-slice instance (their co-execution performs well under MPS when provisioned with 70-30% and 60-40% of the SMs, respectively), but such a configuration is not feasible due to technology limitations. However, in these cases, it could be argued that co-locating applications is not as critical, as the problem arises from their already high resource utilization (see Table 2). This means MIG partitioning lacks the granularity and flexibility to accommodate such fine-grained adjustments.
@@
It is worth noting several nuances regarding this issue. First, we have illustrated this on an A30 GPU because its partitioning space is more limited: it only has 4 slices, 3 instance sizes (1, 2, and 4), and 5 possible partitions, making it unfeasible, for example, to assign 3 or 4 slices to the application with higher memory demand and 2 to the other. On bigger GPUs such as the A100, H100, or B200, which feature higher throughput and greater MIG granularity (7 slices, 5 possible instance sizes, and up to 19 partitions), this issue becomes less relevant because it is, a priori, easier for applications to adapt well to one of the available device partition proportions. Second, MIG includes Compute Instances (see Section 2.4), which allow memory to be shared among concurrent applications within the same instance without sharing the SMs. With these instances, it is possible to obtain results similar to those observed with MPS, although in general cases, this implies predicting and controlling interference to avoid contention problems, which is not always easy, as some scheduling proposals have noted [Xiao2018, Zhang2024]. Finally, the ideal solution would be a MIG-like, but asymmetric, isolation: for example, being able to form an instance with 2 compute slices and 4 memory slices, all of them fully isolated. However, that seems infeasible or technically very difficult, at least in the short term; currently, MIG uses isolation at the GPC level, but it would likely require the ability to isolate resources at a lower, decoupled level.